OpenStudy (anonymous):

Part 1: Provide the solution to the equation 9 = square root of 3 (4x + 6). (3 points) Part 2: Using complete sentences, explain the strategy used in solving this equation the way you did. (3 points)

OpenStudy (anonymous):

\[9 = \sqrt{3}^{4x+6}\] I already got my solution which is x=-1/2, which is part 1. I only need part 2 which i don't know how to explain

OpenStudy (tkhunny):

This may be helpful language: Exploit the BASE portion of each side. \(9 = 3^{2}\) and \(\sqrt{3} = 3^{1/2}\) Rewriting, we have 2 = (1/2)*(4x+6) = 2x+3